question1-ritu is 5 years younger than her brother rohit . Their father is 25 years older than ritu. 4 yeras ago the father was twice the sum of ages of rohit and ritu them .find their ptresent age of each.
Answer by lwsshak3(11628)

question1-ritu is 5 years younger than her brother rohit . Their father is 25 years older than ritu. 4 years ago the father was twice the sum of ages of rohit and ritu then .find their present age of each.
***
let x=ritu's present age
x+5=rohit's present age
25+x=father's present age
..
4 years ago:
x-4= ritu's age
x+5-4=x+1=rohit's age
25+x-4=21+x=father's age
21+x=2(x-4+x+1)=2(2x-3)
21+x=4x-6
3x=27
x=9
x+5=14
25+x=34
ritu's present age=9
rohit's present age=14
father's present age=34